Find the integral with respect to
x:∫cosxsinxdx
1
Expert's answer
2014-09-11T13:53:15-0400
∫cosxsinxdx=(1/2)*∫sin(2x)dx=-(1/2)*cos(2x)/2+C=-(1/4)cos(2x)+C, where C is an arbitrary real constant.